a car at rest initially accelerates uniformly to a speed 108 kmph in 15 s on a straight road. find the distance travelled by the car during this time interval

Grade:

1 Answers

Monil Shah
11 years ago

For distance first find the acceleration using v=u+ at and then substitute in s=half*a*t*t

Answer is 225 m